OpenStudy (anonymous):

If (ax)/2 = bx-1 is solved correctly for x, then the result is ___. (Non-zero divisors.) a. -2/(a-b) b. -1/(2a-b) c. 2/(a+b) d. 2/(2b-a) e. not given Can you also explain how you got it? Thank you :)

OpenStudy (zzr0ck3r):

(ax)/2= bx-1 muliply both sides by two ax = 2bx-2 combine like terms ax-2bx=-2 factor x out x(a-2b)=-2 solve for x x= -2/(a-2b)

OpenStudy (zzr0ck3r):

never mind multiply top and bottom by negative 1 2/(2b-a)

OpenStudy (anonymous):

Why would you multiply by -1?

OpenStudy (lgbasallote):

minor correction.. the transformation from ax = 2bx - 2 into ax - 2bx = -2 is NOT combining like terms, but subtracting 2bx from both sides.. nothing serious..

OpenStudy (zzr0ck3r):

right!

OpenStudy (zzr0ck3r):

I multiply by negative 1 to get it to look like the form in the given answer.

OpenStudy (anonymous):

Shouldn't you multiply -1 to x as well?

OpenStudy (anonymous):

Oh! I got it now. Thank you very much. I appreciate it.

OpenStudy (zzr0ck3r):

yeah that is a little tricky, but since i did it to the top and the bottom, it does not change the ratio, so we would not change the LHS either in other words no we dont multiply x by -1

OpenStudy (anonymous):

That makes sense. I'll try to remember that. Thank you again.